OpenStudy (anonymous):

If t1 = 5 and tn = tn-1 + 2, find the first 4 terms of the sequence.

OpenStudy (lgbasallote):

i strongly advise you to draw that...or retype it using equation button

OpenStudy (anonymous):

Here, \[t_n = t_{n-2} + 2\] \[t_2 = t_1 + 2\] \[t_2 = 7\]

OpenStudy (anonymous):

Like wise Add 2 to the previous term to find the next term..

OpenStudy (anonymous):

what thats how the question is how am i supposed to retype it????

OpenStudy (anonymous):

Tell me when you add 2 to 5 what do you get??

OpenStudy (anonymous):

7

OpenStudy (anonymous):

\(t_n = t_{n-1} + 2\)

OpenStudy (anonymous):

This is your second term.. Keep on adding 2 and find the other two terms,,

OpenStudy (anonymous):

wat other terms?? thats what im not understanding

OpenStudy (anonymous):

You have to find third and fourth term too... what is 7 + 2?? what is 7 + 2 + 2 = ??

OpenStudy (anonymous):

So that will be your 4 terms...

OpenStudy (anonymous):

9

OpenStudy (anonymous):

okaii thank yu now i understnad

OpenStudy (anonymous):

Welcome...